### v1

@@ 2.1 Complexity analysis
We first present some preliminary analysis of our method. The following lemma gives the optimality conditions for the trust region subproblem (2.3), which can be directly obtained from Theorem 4.1 in Jorge1999NumericalO and will not be repeated here.
@@
Lemma 2.4 (Yao2024TwoTR )
Suppose that Assumption 2.1 holds. For any given $\epsilon_{1}>0$ and $\epsilon_{2}>0$, choose
$\eta_{y}=\frac{2}{\ell+\mu},~{}\begin{cases}N_{1}\geqslant\kappa\log\frac{\|y_{-1}-y^{*}(x_{0})\|}{A},\\ N_{t}\geqslant\kappa\log\frac{A+\kappa\|s_{t-1}\|}{A},\hskip 28.45274ptt\geqslant 2,\end{cases}$ (2.8)
where $A=\min\{\epsilon_{1}/l,\epsilon_{2}/L_{H}\}$, then $\|\nabla P(x_{t})-g_{t}\|\leqslant\epsilon_{1}$ and $\|\nabla^{2}P(x_{t})-H_{t}\|\leqslant\epsilon_{2}$.
@@
$\sigma=\frac{\sqrt{L_{2}}}{2},~{}r=\frac{1}{4\sqrt{L_{2}}},~{}\epsilon_{1}=\min\{\frac{1}{96},\frac{\sqrt{L_{2}}}{16L_{1}}\}\epsilon^{3/2},~{}\epsilon_{2}=\frac{\sqrt{L_{2}}}{12}\epsilon^{1/2},$ (2.9)
and choose $N_{t}$ in (2.8). When $\|g_{t}\|\geqslant\epsilon$, if $\lambda_{t}>0$, we have
@@
$g_{t}^{\top}s_{t}+\frac{1}{2}s_{t}^{\top}H_{t}s_{t}=$ $-\frac{1}{2}s_{t}^{\top}(H_{t}+\sigma\|g_{t}\|^{1/2}+\lambda_{t}I)s_{t}-\frac{1}{2}s_{t}^{\top}(\sigma\|g_{t}\|^{1/2}+\lambda_{t}I)s_{t}$
$\leqslant$ $-\frac{1}{2}s_{t}^{\top}(\sigma\|g_{t}\|^{1/2}+\lambda_{t}I)s_{t}.$ (2.13)
@@
$P(x_{t+1})\leqslant P(x_{t})-\frac{1}{2}s_{t}^{\top}(\sigma\|g_{t}\|^{1/2}+\lambda_{t}I)s_{t}+\epsilon_{1}\|s_{t}\|+\frac{\epsilon_{2}}{2}\|s_{t}\|^{2}+\frac{L_{2}}{6}\|s_{t}\|^{3}.$ (2.14)
@@
$\leqslant$ $-(\frac{\sigma}{2}-\frac{L_{2}r}{6})\|g_{t}\|^{1/2}\|s_{t}\|^{2}+\min\{\frac{1}{96},\frac{\sqrt{L_{2}}}{16L_{1}}\}\epsilon^{3/2}\|s_{t}\|+\frac{\sqrt{L_{2}}}{24}\epsilon^{1/2}\|s_{t}\|^{2}$
$\leqslant$ $-(\frac{\sqrt{L_{2}}}{6}-\min\{\frac{1}{96},\frac{\sqrt{L_{2}}}{16L_{1}}\}\frac{\epsilon^{3/2}}{\|g_{t}\|^{1/2}\|s_{t}\|})\|g_{t}\|^{1/2}\|s_{t}\|^{2}.$ (2.18)
@@
$~{}\epsilon^{2/3}\leqslant\epsilon^{1/2}\|g_{t}\|\leqslant(L_{1}+\sigma)\|g_{t}\|^{1/2}\|s_{t}\|.$ (2.19)
@@
$\leqslant$ $~{}(2\frac{\epsilon_{1}}{\|g_{t}\|}+\frac{L_{2}}{2}r^{2}+\frac{\epsilon_{2}}{\|g_{t}\|^{1/2}}r+\sigma r)\|g_{t}\|.$ (2.20)
@@
$\mathcal{F}_{\tilde{T}(\epsilon)}=\big{\{}t<\tilde{T}(\epsilon):P(x_{t+1})-P(x_{t})\leqslant-\frac{1}{128\sqrt{L_{2}}}\max\{\|g_{t}\|^{3/2},\epsilon^{3/2}\}\big{\}},$ (2.25)
@@
According to Lemmas 2.5 and 2.6, each iteration $t$ belongs to one of the two sets $\mathcal{F}_{\tilde{T}(\epsilon)}$ and $\mathcal{G}_{\tilde{T}(\epsilon)}$. Therefore, our aim is to establish an upper bound for the cardinality of sets $\mathcal{F}_{\tilde{T}(\epsilon)}$ and $\mathcal{G}_{\tilde{T}(\epsilon)}$.
@@
According to Lemmas 2.5 and 2.6, for any $t\geqslant 1$, we have
@@
$\leqslant$ $\sum_{t=1}^{\tilde{T}(\epsilon)}(P(x_{t})-P(x_{t+1}))\leqslant P(x_{0})-P^{*},$
@@
$|\mathcal{G}_{\tilde{T}(\epsilon)}|\leqslant\lceil\log(G/\epsilon)/\log(3)\rceil|\mathcal{F}_{\tilde{T}(\epsilon)}|,$ (2.30)
@@
$\|g_{t+1}\|\leqslant G=\max\big{\{}2+\frac{L_{1}}{4\sqrt{L_{2}}}C^{1/2}+C,2+\frac{L_{1}}{4\sqrt{L_{2}}}\big{\}}.$
@@
Next we give an upper bound on the index set $|\mathcal{G}_{\tilde{T}(\epsilon)}|$. Denote the maximum number of consecutive iterates in set $\mathcal{G}_{\tilde{T}(\epsilon)}$ as $n$. We prove $n\leqslant\lceil\log(G/\epsilon)/\log(3)\rceil$ by contradiction. Assume $n>\log(G/\epsilon)/\log(3)$. From $t\in\mathcal{G}_{\tilde{T}(\epsilon)}$ and $\|g_{t}\|\leqslant G$, we know that
$\|g_{t+n}\|\leqslant(\frac{1}{3})^{n}\|g_{t}\|\leqslant(\frac{1}{3})^{n}G<\epsilon,$
Then according to Lemma 2.6, $t+n\in\mathcal{F}_{\tilde{T}(\epsilon)}$, which contradicts the definition of $n$. Therefore, $n\leqslant\lceil\log(G/\epsilon)/\log(3)\rceil$ and the inequality (2.30) follows directly.
@@
$T(\epsilon)\leqslant 128(\lceil\log(G/\epsilon)/\log(3)\rceil+1)\sqrt{L_{2}}(P(x_{0})-P^{*})\epsilon^{-3/2}+1,$ (2.32)
where $G$ is defined as in Lemma 2.8.
The conclusion of the theorem can be directly proved by Lemmas 2.7 and 2.8, and the details are not elaborated here.
Remark 1

### v2

@@ 2.1 Complexity analysis
We first present some preliminary analysis of our method. The following lemma gives the optimality conditions for the trust-region subproblem (2.3), which can be directly obtained from Theorem 4.1 in Jorge1999NumericalO and will not be repeated here.
@@
Lemma 2.4
Suppose that Assumption 2.1 holds. For any given $\epsilon_{1}>0$ and $\epsilon_{2}>0$, set
$\eta_{y}=\frac{1}{\ell},~{}\theta=\frac{\sqrt{\kappa}-1}{\sqrt{\kappa}+1},~{}\begin{cases}N_{0}\geqslant 2\sqrt{\kappa}\ln\frac{\sqrt{\kappa+1}\|y_{-1}-y^{*}(x_{0})\|}{A},\\ N_{t}\geqslant 2\sqrt{\kappa}\ln\frac{\sqrt{\kappa+1}(A+\kappa\|s_{t-1}\|)}{A},t\geqslant 1,\end{cases}$ (2.8)
where $A=\min\{\epsilon_{1}/\ell,\epsilon_{2}/(2L_{H})\}$, then $\|\nabla P(x_{t})-g_{t}\|\leqslant\epsilon_{1}$ and $\|\nabla^{2}P(x_{t})-H_{t}\|\leqslant\epsilon_{2}$.
Proof
We first use induction to show that
$\|y_{t}-y^{*}(x_{t})\|\leqslant A$ (2.9)
for all $t\geqslant 0$. Note that $y_{t}$ is obtained by applying $N_{t}$ Nesterov’s accelerated gradient ascent steps starting from $y_{t-1}$. By Assumption 2.1 and (2.2.23) in Nesterov2018Lectures , we have
$\|y_{t}-y^{*}(x_{t})\|\leqslant$ $~{}(\sqrt{\kappa+1})e^{-N_{t}/(2\sqrt{\kappa})}\|y_{t-1}-y^{*}(x_{t})\|.$ (2.10)
By further combining with (2.8), (2.9) holds directly for $t=0$. Suppose that (2.9) holds for $t$, by (2.10), Lemma 2.1 and (2.8), we then obtain that
$\|y_{t+1}-y^{*}(x_{t+1})\|\leqslant$ $~{}(\sqrt{\kappa+1})e^{-N_{t+1}/(2\sqrt{\kappa})}\|y_{t}-y^{*}(x_{t+1})\|$
$\leqslant$ $~{}(\sqrt{\kappa+1})e^{-N_{t+1}/(2\sqrt{\kappa})}(\|y_{t}-y^{*}(x_{t})\|+\|y^{*}(x_{t+1})-y^{*}(x_{t})\|)$
$\leqslant$ $~{}(\sqrt{\kappa+1})e^{-N_{t+1}/(2\sqrt{\kappa})}(A+\kappa\|x_{t+1}-x_{t}\|)\leqslant A,$
which completes the proof of (2.9). Then, by Assumption 2.1, Lemma 2.1, Lemma 2.2 and (2.9), we get
$\|\nabla P(x_{t})-g_{t}\|$ $=\|\nabla_{x}f(x_{t},y^{*}(x_{t}))-\nabla_{x}f(x_{t},y_{t})\|\leqslant l\|y_{t}-y^{*}(x_{t})\|\leqslant\epsilon_{1},$
$\|\nabla^{2}P(x_{t})-H_{t}\|$ $=\|H(x_{t},y^{*}(x_{t}))-H(x_{t},y_{t})\|\leqslant L_{H}\|y_{t}-y^{*}(x_{t})\|\leqslant\epsilon_{2}.$
The proof is then completed.
@@
$\sigma=\frac{\sqrt{L_{2}}}{2},~{}r=\frac{1}{4\sqrt{L_{2}}},~{}\epsilon_{1}=\min\left\{\frac{1}{96},\frac{\sqrt{L_{2}}}{16L_{1}}\right\}\epsilon^{3/2},~{}\epsilon_{2}=\frac{\sqrt{L_{2}}}{12}\epsilon^{1/2},$ (2.11)
and choose $\eta_{y},\theta,N_{t}$ in (2.8). When $\|g_{t}\|\geqslant\epsilon$, if $\lambda_{t}>0$, we have
@@
$g_{t}^{\top}s_{t}+\frac{1}{2}s_{t}^{\top}H_{t}s_{t}=$ $-\frac{1}{2}s_{t}^{\top}(H_{t}+\sigma\|g_{t}\|^{1/2}I+\lambda_{t}I)s_{t}-\frac{1}{2}s_{t}^{\top}(\sigma\|g_{t}\|^{1/2}I+\lambda_{t}I)s_{t}$
$\leqslant$ $-\frac{1}{2}s_{t}^{\top}(\sigma\|g_{t}\|^{1/2}I+\lambda_{t}I)s_{t}.$ (2.15)
@@
$P(x_{t+1})\leqslant P(x_{t})-\frac{1}{2}s_{t}^{\top}(\sigma\|g_{t}\|^{1/2}I+\lambda_{t}I)s_{t}+\epsilon_{1}\|s_{t}\|+\frac{\epsilon_{2}}{2}\|s_{t}\|^{2}+\frac{L_{2}}{6}\|s_{t}\|^{3}.$ (2.16)
@@
$\leqslant$ $-\bigg{(}\frac{\sigma}{2}-\frac{L_{2}r}{6}\bigg{)}\|g_{t}\|^{1/2}\|s_{t}\|^{2}+\min\bigg{\{}\frac{1}{96},\frac{\sqrt{L_{2}}}{16L_{1}}\bigg{\}}\epsilon^{3/2}\|s_{t}\|+\frac{\sqrt{L_{2}}}{24}\epsilon^{1/2}\|s_{t}\|^{2}$
$\leqslant$ $-\left(\frac{\sqrt{L_{2}}}{6}-\min\bigg{\{}\frac{1}{96},\frac{\sqrt{L_{2}}}{16L_{1}}\bigg{\}}\frac{\epsilon^{3/2}}{\|g_{t}\|^{1/2}\|s_{t}\|}\right)\|g_{t}\|^{1/2}\|s_{t}\|^{2}.$ (2.20)
@@
$~{}\epsilon^{3/2}\leqslant\epsilon^{1/2}\|g_{t}\|\leqslant(L_{1}+\sigma)\|g_{t}\|^{1/2}\|s_{t}\|.$ (2.21)
@@
$\leqslant$ $~{}\left(2\frac{\epsilon_{1}}{\|g_{t}\|}+\frac{L_{2}}{2}r^{2}+\frac{\epsilon_{2}}{\|g_{t}\|^{1/2}}r+\sigma r\right)\|g_{t}\|.$ (2.22)
@@
$\mathcal{F}_{\tilde{T}(\epsilon)}=\bigg{\{}t<\tilde{T}(\epsilon):P(x_{t+1})-P(x_{t})\leqslant-\frac{1}{128\sqrt{L_{2}}}\max\{\|g_{t}\|^{3/2},\epsilon^{3/2}\}\bigg{\}},$ (2.27)
@@
According to Lemmas 2.5 and 2.6, each iteration $t$ belongs to at least one of the two sets $\mathcal{F}_{\tilde{T}(\epsilon)}$ and $\mathcal{G}_{\tilde{T}(\epsilon)}$. Therefore, our aim is to establish an upper bound for the cardinality of sets $\mathcal{F}_{\tilde{T}(\epsilon)}$ and $\mathcal{G}_{\tilde{T}(\epsilon)}$.
@@
According to Lemmas 2.5 and 2.6, for any $t\geqslant 0$, we have
@@
$\leqslant$ $\sum_{t=0}^{\tilde{T}(\epsilon)-1}(P(x_{t})-P(x_{t+1}))\leqslant P(x_{0})-P^{*},$
@@
$|\mathcal{G}_{\tilde{T}(\epsilon)}|\leqslant\lceil\ln(G/\epsilon)/\ln 3\rceil|\mathcal{F}_{\tilde{T}(\epsilon)}|,$ (2.32)
@@
$\|g_{t+1}\|\leqslant G=\max\left\{2+\frac{L_{1}}{4\sqrt{L_{2}}}C^{1/2}+C,2+\frac{L_{1}}{4\sqrt{L_{2}}}\right\}.$
@@
Next we give an upper bound on the index set $|\mathcal{G}_{\tilde{T}(\epsilon)}|$. Denote the maximum number of consecutive iterates in set $\mathcal{G}_{\tilde{T}(\epsilon)}$ as $n$. We prove $n\leqslant\lceil\ln(G/\epsilon)/\ln(3)\rceil$ by contradiction. Assume $n>\ln(G/\epsilon)/\ln(3)$. From $t\in\mathcal{G}_{\tilde{T}(\epsilon)}$ and $\|g_{t}\|\leqslant G$, we know that
$\|g_{t+n}\|\leqslant\big{(}\frac{1}{3}\big{)}^{n}\|g_{t}\|\leqslant\big{(}\frac{1}{3}\big{)}^{n}G<\epsilon,$
Then according to Lemma 2.6, $t+n\in\mathcal{F}_{\tilde{T}(\epsilon)}$, which contradicts the definition of $n$. Therefore, $n\leqslant\lceil\ln(G/\epsilon)/\ln(3)\rceil$ and the inequality (2.32) follows directly. This completes the proof.
@@
$T(\epsilon)\leqslant$ $~{}128(\lceil\ln(G/\epsilon)/\ln 3\rceil+1)\sqrt{L_{2}}(P(x_{0})-P^{*})\epsilon^{-3/2}+1,$
$\sum_{t=0}^{T(\epsilon)}N_{t}\leqslant 2\sqrt{\kappa}$ $~{}\ln\frac{\sqrt{\kappa+1}\|y_{-1}-y^{*}(x_{0})\|}{A}+2\sqrt{\kappa}T(\epsilon)\ln\frac{\sqrt{\kappa+1}(A+\kappa r\sqrt{G})}{A},$
where $G$ is defined as in Lemma 2.8 and $A$ is defined as in Lemma 2.4.
Proof
The upper bound of $T(\epsilon)$ can be directly proved by Lemmas 2.7 and 2.8, and the specific details will not be repeated here. According to Lemma 2.4, the total number of gradient ascent steps required satisfies
$\sum_{t=0}^{T(\epsilon)}N_{t}=$ $~{}N_{0}+\sum_{t=1}^{T(\epsilon)}2\sqrt{\kappa}\ln\frac{\sqrt{\kappa+1}(A+\kappa\|s_{t-1}\|)}{A}$
$=$ $~{}N_{0}+2\sqrt{\kappa}\ln\bigg{(}\prod_{t=1}^{T(\epsilon)}\frac{\sqrt{\kappa+1}(A+\kappa\|s_{t-1}\|)}{A}\bigg{)}$
$\leqslant$ $~{}N_{0}+2\sqrt{\kappa}T(\epsilon)\ln\bigg{(}\frac{1}{T(\epsilon)}\sum_{t=1}^{T(\epsilon)}\frac{\sqrt{\kappa+1}(A+\kappa\|s_{t-1}\|)}{A}\bigg{)}.$
By further combining with $\|s_{t}\|\leqslant r\max\{\epsilon^{1/2},\|g_{t}\|^{1/2}\}$ and $\|g_{t}\|\leqslant G$ in Lemma 2.8, we conclude that
$\sum_{t=0}^{T(\epsilon)}N_{t}\leqslant 2\sqrt{\kappa}\ln\frac{\sqrt{\kappa+1}\|y_{-1}-y^{*}(x_{0})\|}{A}+2\sqrt{\kappa}T(\epsilon)\ln\frac{\sqrt{\kappa+1}(A+\kappa r\sqrt{G})}{A},$
which completes the proof.
Remark 2.1